OpenStudy (moongazer):

The speed of a car is 80 miles per hour. Find the radius of the car wheels in inches if the car makes 300 rev/min.

OpenStudy (moongazer):

I got r=44.82 inches is it correct?

OpenStudy (amistre64):

how many revolutions per hour does it make?

OpenStudy (moongazer):

5 rev per hour

OpenStudy (amistre64):

really? it does 300 revs in one minute; how can it do way less revs in one hour?

OpenStudy (amistre64):

300*60 ...

OpenStudy (moongazer):

sorry for that I interchanged it :) 18000 per hour

OpenStudy (amistre64):

18000 revs in one hour put to the ratio of 80 miles 18000/80 is our next step i believe

OpenStudy (amistre64):

of if we knew how many inches 80 miles was we could go more directly to an answer ....

OpenStudy (moongazer):

why did you made it into ratio?

OpenStudy (amistre64):

becasue that is what it is: 18000 revs = 80 miles i spose since we want to know how far 1 rev takes us its better to make the rev side equal to 1 tho divide both sides by 18000

OpenStudy (amistre64):

1 rev = the circumference of the wheel

OpenStudy (moongazer):

sorry I got confused :)

OpenStudy (amistre64):

if we convert 80 miles to inches; then we know how many inches 1 rev is; so we would know how many inches the circumference of the wheel is knowing circunference is equal to 2pi*radius; we solve for the radius r = C/2pi

OpenStudy (moongazer):

Sorry, but I am really confused with your method. :) I did angular and linear speed with this one.

OpenStudy (amistre64):

google says: 1mile=63360inches multiply both sides by 80 we get 80 miles = 5 068 800 inches 18000 revs = 5 068 800 inches divide both sides by 18000 to get 1 revolution 5 068 800/ 18000 = 281.6 281.6 inches is our circumference therefore, using what i posted up above; radius = 281.6/(2pi)

OpenStudy (amistre64):

i did the same, i converted angular (rpm) into linear (miles, inches)

OpenStudy (amistre64):

of course i get about a wheel that is about 7 feet in diameter lol

OpenStudy (moongazer):

brb. i'll try to understand your method. it looks good :) still, I think we got the same answer but different method of getting it.

OpenStudy (amistre64):

44.82, yes

OpenStudy (moongazer):

and I think your method is faster :)